OpenStudy (anonymous):

How to solve this equation 441=40x+x^2 ????

OpenStudy (anonymous):

\[441=40x+x ^{2}\]

OpenStudy (anonymous):

start with \[x^2+40x-441=0\] if you want to factor otherwise it is probably easiest to complete the square the problem was cooked up to factor because \[x^2+40x-441=0\] factors as \[(x-9)(x+49)=0\] but it is not clear how you would know that

OpenStudy (anonymous):

however, once you know it factors as \[(x-9)(x+49)=0\] it is easy to solve \[x-9=0\iff x=9\]\[x-49=x\iff x=-49\]

OpenStudy (anonymous):

I don't understand, can you explain what your doing please?

OpenStudy (anonymous):

if you did not see that it factors, you can always complete the square starting with \[x^2+40x=441\] you get \[(x+20)^2=441+20^2=841\] and so \[x+20=\pm\sqrt{841}=\pm29\] etc

OpenStudy (anonymous):

ok lets go slow first the factor method you had \[441=40x+x ^{2}\]

OpenStudy (anonymous):

subtracting \(441\) from both sides give \[x^2+40x-441=0\] ok so far?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

then you need to factor this it factors as \((x-9)(x+29)=0\) this is because \(-9\times 49=-441\) and \(49-9=40\)

OpenStudy (anonymous):

finding those factors is some magic, that is why i like completing the square more but assuming we know that \[x^2+40x-441=0\] is the same as \[(x-9)(x+49)=0\] then we can solve easily

OpenStudy (anonymous):

set each factor equal to zero and solve for \(x\) because if the product is zero, then one of the factors has to be zero i wrote above \(x-9=0\iff x=9\) right?

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

Got it!

OpenStudy (anonymous):

and also \(x+49=0\iff x=-49\)

OpenStudy (anonymous):

Thanks you so much! Your awesome!!!!!

OpenStudy (anonymous):

the hard part is factoring, i just said it, didn't show how to do it, but if it is clear, then so much the better yw